OpenStudy (anonymous):

Find an explicit rule for the nth term of the sequence. -5, -25, -125, -625, ...

OpenStudy (welshfella):

can you see the pattern - what value multiplies each term to get the next one?

OpenStudy (welshfella):

-5 * x = -25 -25 * x = -125 what is x?

OpenStudy (anonymous):

Well it looks like it would be a positive number since it's constant negative

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

x = 5

OpenStudy (welshfella):

thats correct

OpenStudy (anonymous):

Here are the options, forgot to post them an = -5 • 5^n an = 5 • -5^(n + 1) an = 5 • -5^n an = -5 • 5^(n - 1)

OpenStudy (welshfella):

this is a geometric sequence with common ratio r = 5 if we represent the first term by a1 then first term = a1 2nd = a1 r 3rd = a1 r^2 4th = a1 r^3 so nth term = ?

OpenStudy (anonymous):

a1r^8

OpenStudy (anonymous):

What is nth, 9th?

OpenStudy (welshfella):

No thats the 9th term

OpenStudy (welshfella):

instead of 3,4,5th term we write n and since the exponent is one less than the n we write the exponent as n-1

OpenStudy (welshfella):

when i said 3,4 5th and didnt mean just these - it applies to all terms

OpenStudy (anonymous):

Oh okay that makes sense

OpenStudy (welshfella):